# Soft-CAM: Making black box models self-explainable for high-stakes decisions

### v1

@@ Abstract
Convolutional neural networks (CNNs) are widely used for high-stakes applications like medicine, often surpassing human performance. However, most explanation methods rely on post-hoc attribution, approximating the decision-making process of already trained black-box models. These methods are often sensitive, unreliable, and fail to reflect true model reasoning, limiting their trustworthiness in critical applications. In this work, we introduce SoftCAM, a straightforward yet effective approach that makes standard CNN architectures inherently interpretable. By removing the global average pooling layer and replacing the fully connected classification layer with a convolution-based class evidence layer, SoftCAM preserves spatial information and produces explicit class activation maps that form the basis of the model’s predictions. Evaluated on three medical datasets, SoftCAM maintains classification performance while significantly improving both the qualitative and quantitative explanation compared to existing post-hoc methods. Our results demonstrate that CNNs can be inherently interpretable without compromising performance, advancing the development of self-explainable deep learning for high-stakes decision-making.
@@ 1 Introduction
Convolutional Neural Networks (CNNs) have revolutionized computer vision by efficiently capturing local patterns, reducing parameters, and accelerating convergence, enabling superior performance in tasks like image recognition and object detection [35, 40]. However, their lack of interpretability limits adoption in high-stakes fields like medicine, where transparency and trust are crucial. To explain CNNs, numerous saliency-based methods have been proposed, including class activation maps (CAM) [61] and their variants [14, 47, 56, 59], gradient-based techniques [48, 50, 51, 53], and even perturbation- or occlusion-based methods [22, 56, 59]. These methods have been widely adopted to explain the decisions of trained black-box models.
Such saliency map-based techniques offer explanations for CNN classifiers that claim to highlight regions in the input image most relevant to the model’s prediction. These explanations are generated post-hoc, typically after a model is trained [10, 27]. Studies have shown significant limitations in their effectiveness, especially in clinical settings [5]. Post-hoc saliency methods often lack faithfulness, reliability, and consistency, resulting in explanations that may not accurately reflect the model’s decision-making process [3, 58]. Moreover, they struggle to localize relevant regions in medical imaging [4], where the limited availability of ground-truth annotations makes it difficult to assess their trustworthiness. To overcome these challenges, inherently and/or self-explainable models have been introduced [45], designed explicitly to provide interpretable insights by incorporating explanations within their architecture [11, 13, 16, 18, 52]. These models generate more trustworthy and faithful explanations that align closely with the model’s actual reasoning [28]. However, self-explainable models generally use specific architectures [13, 16, 52], which limits their applicability and generalization to widely used CNN architectures.
Motivated by these challenges, we propose SoftCAM, a straightforward generalization of Class Activation Maps (CAM) that uses a convolution-based classifier to transform any black-box CNN into a self-explainable model. By removing the final pooling layer and replacing the fully connected classification layer with 1x1 convolutions, SoftCAM turns classical CNNs into fully convolutional networks, generating class-specific evidence maps that are directly used for predictions. Our contributions are:
  • We introduced SoftCAM, a simple modification to CNNs that enables self-explainability, and experimentally demonstrate that the resulting model preserves classification performance across three clinically relevant medical datasets spanning different imaging modalities.
  • We showed that regularizing evidence maps using ElasticNet, a regularizer combining both ridge and lasso penalties, enhances the model’s explanations.
  • We evaluated five widely used traditional CAM-based post-hoc explanation methods, showing that SoftCAM most often outperforms them across a broad range of explainability metrics and across all three considered medical imaging datasets and modalities.
2 Related work
Deep neural networks (DNNs) are widely used in a variety of fields. However, regulatory frameworks such as the European AI Act require AI-based decisions to be explainable to ensure fairness, transparency, and accountability, allowing users to understand their decision-making process [2, 41]. Explainability can help verify and even improve performance by detecting shortcuts and identifying clinically relevant features, ultimately fostering greater trust in decision-making, especially in fields like healthcare [17].
Existing explainable AI (XAI) methods for image analysis can be broadly categorized into attribution-based and non-attribution-based approaches [10, 27, 57]. Attribution-based methods explain “where” important features exist in the input by generating saliency or heatmaps that assign importance scores to individual pixels or regions, helping visualize their contribution to the model’s decision. These include perturbation-based methods [29] and class activation maps [26], which consist of both gradient-based [47, 51, 53] and gradient-free approaches [44, 56, 61]. In contrast, non-attribution-based approaches explain “why” a decision was made without relying on importance scores, instead using techniques such as concept-based (ACE [23], TCAV [32], CBM [33]), prototype-based [15, 16], or counterfactual-based [7, 12, 24, 52] methods to analyze model behavior from different perspectives. These approaches also differ in how explanations are obtained and which architectures they can be applied to. Attribution-based methods typically provide post-hoc, local explanations by offering input-specific insights into black-box CNN models after training. In contrast, non-attribution-based methods are generally inherently interpretable by design, promoting transparency and enabling global understanding of the model’s decision-making process across the entire dataset [45]. However, in some cases, self-explainable models may be less effective for complex tasks, highlighting a tradeoff between interpretability and performance, where increasing transparency may sometimes come at the cost of classification performance [57].
Our method, SoftCAM, relies on an explicit class-evidence layer based on convolutional operations for classification, offering class-specific, attribution-based explanations while remaining inherently interpretable, unlike post-hoc approaches. Moreover, it maintains predictive performance comparable to its corresponding non-interpretable black-box models. Closely related work includes [6] and [18]. In [6], a dual-branch approach is used, where one branch leverages a traditional black-box CNN model (ResNet-18) for classification, and the second branch uses weight-sharing for post-hoc explanations, requiring two forward passes for inference. In the second branch for post-hoc explanation, the global average pooling layer (GAP) is removed, and the linear classifier is replaced by convolutional layers that share weights during inference to generate class-specific activation maps. In contrast, SoftCAM is trained end-to-end, providing both predictions and explanations in a single forward pass, eliminating the need for additional computational overhead or weight sharing. Furthermore, while [18] uses explicit class-evidence maps to enhance the explanation of a self-explainable bag-of-local-feature model (BagNet [13]), our method transforms black-box models into self-explainable models. We evaluated our approach on a range of medical datasets, comparing the resulting explanations to various post-hoc attribution-based methods, including both gradient-free and gradient-based techniques.
@@ 3 Method
Given an input image $\mathbf{X}\in\mathbb{R}^{H_{X}\times W_{X}\times C_{X}}$ with height $H_{X}$, width $W_{X}$, and the number of channels $C_{X}$, consider a CNN network $f_{\theta}$ that maps $\mathbf{X}$ to a probability distribution $\mathbf{\hat{y}}=f_{\theta}(\mathbf{X})\in\mathbb{R}^{C}$, where $C$ is the number of classes, and $y^{c}\in\mathbf{y}$ represents the predicted probability for class $c$. The network consists of a feature extractor $g_{\phi}$, and a classifier layer $h_{\psi}$, with learnable parameters $\phi$ and $\psi$. The feature extractor generates a feature map $\mathbf{Z}=g_{\phi}(\mathbf{X})\in\mathbb{R}^{N\times M\times D}$, where $N\times M$ denotes the spatial size and $D$ is the feature dimension (e.g., $D=2048$ for most ResNet variants). The classifier then predicts the final output based on the extracted features. Let $\mathcal{A}=\{\mathbf{A}_{k}\}_{k=1}^{D}$, the set of activation maps obtained from the feature extractor, where $A_{k}$ is the activation of the $k$-th neuron. Let, $S^{c}_{\text{Map}}\in\mathbb{R}^{N\times M}$ be the 2D saliency map, providing a visual explanation of the model’s prediction for class $c$. This paper explores how to train self-explainable CNNs to simultaneously generate both the prediction $y^{c}$ and its corresponding explanation $S^{c}_{\text{Map}}$.
Traditional CNN architectures employ a GAP layer to reduce the feature map to $1\times D$, followed by a classification module consisting of one or more linear fully connected layers (FCLs) to generate the final prediction. Post-hoc methods are then typically used to explain the model’s decision.
@@ 3.1 CAM-based methods
Class Activation Maps (CAMs) [61] are closely related to our approach, offering local visual explanations of CNN predictions by generating saliency maps for individual inputs. CAM achieves this by linearly combining the feature maps from the final convolutional layers with importance coefficients from the FCL classifier, thereby producing class-wise attribution maps as follows:
$S_{\text{CAM}}^{c}(x_{1},x_{2})=\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2}),$ (1)
where $A_{k}(x_{1},x_{2})$ is the activation of neuron $k$ in the feature map at spatial location $(x_{1},x_{2})$, and $w_{k}^{c}$ denotes the importance weight associated with class $c$ for unit $k$ in the fully connected layer.
Originally, CAM was designed for CNNs with GAP and FCL, but has been extended to gradient-based methods using class score gradients to compute importance weights [14, 47, 48]. This extension enabled CAM-based techniques to be applied to a broader range of CNN architectures, particularly those where the GAP layer is followed by multiple FCLs, as seen in models like VGG [49] and InceptionV3 [54]. For example, GradCAM [47] extends the original CAM approach by backpropagating the gradient from a target class to the input layer to highlight the image regions that strongly influence the model’s prediction. GradCAM is formulated as
$S^{c}_{\text{Grad-CAM}}(x_{1},x_{2})=\text{ReLU}\Bigg{(}\sum_{k=1}^{D}w_{k}^{c}A_{k}(x_{1},x_{2})\Bigg{)},$ (2)
where the weight coefficients are computed as $w_{k}^{c}=\frac{1}{N\times M}\sum_{i}^{N}\sum_{j}^{N}\frac{\partial y^{c}}{\partial A_{k}(i,j)}$. Here, $A_{k}(i,j)$ is the activation value at location $(i,j)$ on $A_{k}$, and the rectified linear unit (ReLU) is applied to ensure that the final activation map considers only the features that positively influence class $c$. Following GradCAM, several variations have been proposed, including gradient-based approaches such as SmoothGrad [50], GradCAM++ [14], guided-backpropagation [51], and integrated gradients [53], as well as gradient-free methods like ScoreCAM [56], LayerCAM [30], and OptiCAM [59]. Gradient-based methods primarily differ in how gradients are aggregated to compute importance weights, while gradient-free methods mainly vary in how the weights are computed.
Despite the success of class activation map-based methods in explaining CNN classifiers, including medical applications [8], they have a key limitation: they rely on already trained models and provide post-hoc explanations, which may not accurately reflect the model’s true decision-making process. Additionally, gradient-based methods face inherent challenges such as gradient saturation, where DNN gradients tend to diminish, and false confidence, where the highest activation map weight does not necessarily correspond to the greatest increase in confidence [56]. On the other hand, gradient-free methods are computationally and memory-intensive, as they often require multiple forward passes on perturbed inputs. Finally, CAM-based methods are easy to implement in CNNs with clearly defined spatial feature maps, but face challenges in multi-branch architectures like InceptionV3 due to the complexity of integrating diverse feature maps from parallel convolutional paths.
@@
3.2 Improving CAMs for self-explanability
Motivated by the limitations of post-hoc class activation map-based methods in interpreting CNN classifiers, we introduce SoftCAM (Fig. 1), a straightforward modification of black-box CNN classifiers that makes them self-explainable and inherently interpretable. SoftCAM achieves this by replacing the fully connected classification layer in classical CNNs with an explicit class-evidence convolutional layer, preserving spatial information and providing explanations in a single forward pass, eliminating the need and computational overhead for post-hoc techniques.
We make black-box CNN architectures self-explainable by modifying how predictions are obtained. Any FCL of size $b_{1}\times b_{2}$, where $b_{1}$ and $b_{2}$ denote the number of input and output features, respectively, can be equivalently expressed as a $1\times 1$ convolutional layer with $b_{1}$ input channels and $b_{2}$ output channels [18]. This allows us to replace FCL classifiers in standard CNN architectures with convolutions, removing the GAP layer before classification, while preserving model complexity and spatial localization. The new classifier module $h$ consists of convolutional layers (Fig. 1d) with $C$ convolution kernels of size $1\times 1$ and unit stride, producing class evidence maps (Fig. 1e)
@@ 3.2 Improving CAMs for self-explanability
where $\psi$ is a learnable parameter. Indeed, $h_{\psi}$ can be viewed as an explainable, soft generalization of classical post-hoc attribution methods (Eq. 1, 2), mapping the low-dimensional feature volume $\mathbf{Z}$ into an interpretable, class-wise activation volume $\mathbf{A}$ whose reduced channel dimension corresponds to the number of target classes. Unlike CAM (Eq. 1) and GradCAM (Eq. 2), which generate post-hoc heuristic explanations, our approach leverages the feature map volume from the backbone and applies a parameterized function $h_{\psi}$ to directly produce class activation maps that are used for prediction. In contrast to classical CAM-based methods, the importance weights are not explicitly defined but are implicitly learned and encoded within the classifier’s parameters.
The resulting architecture is a fully convolutional, self-explainable model, where the final predicted probabilities are computed from the evidence map (Fig.1e), without introducing additional parameters:
$\mathbf{\hat{y}}=\text{Softmax}\bigg{(}\mbox{AvgPool}\Big{(}h_{\psi}\big{(}g_{\Phi}(\mathbf{X})\big{)}\Big{)}\bigg{)}\in\mathbb{R}^{1\times C}.$ (4)
Additionally, the class evidence maps $\mathbf{A}$ serve as built-in explanations, directly representing the contribution of individual input regions to the final prediction (Fig. 1g). Replacing linear FCL layers with convolutional operations offers several advantages. Due to the shift-invariance and position-agnostic properties of CNNs, all image regions are weighted equally when forming the final classification (Fig. 1f). As a result, input feature patches with high activations in the evidence maps contribute most significantly and linearly to the prediction. This behavior mirrors that of simple linear models, where each value in the activation map has a direct and interpretable impact on the output.
@@ 3.3 Regularizing SoftCAM for interpretability
By using explicit class-evidence maps, the model can be trained directly with regularization applied to these maps to enhance interpretability. In practice, we apply an ElasticNet regularization [62], which linearly combines the $\ell_{1}$ (Lasso) and $\ell_{2}$ (Ridge) penalties, leading to the following loss function:
$\mathcal{L}(\mathbf{y},\mathbf{\hat{y}})=\mbox{CE}(\mathbf{y},\mathbf{\hat{y}})+\lambda_{1}\sum_{i,j,c}|\mathbf{A}_{c}^{ij}|+\lambda_{2}\sum_{i,j,c}||\mathbf{A}_{c}^{ij}||_{2}.$ (5)
Here, CE denotes the cross-entropy loss, and $\mathbf{y}$ represents the reference class labels. When $\lambda_{2}=0$, the ElasticNet penalty becomes the Lasso penalty, which promotes sparsity in the class evidence maps [18] by removing less informative activations, making it particularly useful for tasks where precision in explanations is crucial. In contrast, when $\lambda_{1}=0$, the ElasticNet reduces to the Ridge penalty, which reduces irrelevant activations without forcing them to zero, which is beneficial when minimizing false negatives is a priority. ElasticNet thus provides a balance between Lasso and Ridge penalties, balancing sparsity and smoothness in the resulting activation maps.
Visualizing explanations.
The evidence map generated by SoftCAM is upsampled to the input resolution for visualization (Fig. 1g). Like most CAM-based methods, such as GradCAM [47], ScoreCAM [56], and LayerCAM [30] that operate on the final convolutional layer, SoftCAM’s explanations are limited by the resolution of the backbone (e.g., 16×16 for VGG-16/ResNet-50 with 512×512 input) due to pooling and striding, leading to lower-resolution saliency maps. However, by introducing the class evidence and classification layer directly atop features, SoftCAM regularizes the evidence map, making it less coarse and thereby enhancing localization. In contrast, gradient-based methods like Integrated Gradients [53] and Guided Backpropagation [51] produce high-resolution saliency maps by computing pixel-level gradients, which may lead to noisy maps, especially when the region of interest spans a broader area, as commonly observed in Chest X-ray images.
Comparison with other approaches.
Unlike post-hoc attribution-based approaches, our method is inherently interpretable from the classification layer and maintains performance comparable to its black-box counterpart, without a significant trade-off, even when regularization is applied to enhance explainability. Compared to [18], our method extends from the concept of interpretable bag-of-local models to general black-box CNN architectures and generalizes the regularization from Lasso to ElasticNet, with extensive evaluations across multiple datasets using a broad range of explanability metrics. Compared to [6], our method is trained end-to-end and does not require post-hoc processing, weight sharing between branches, or an additional forward pass to generate explanations.

### v2

@@ Abstract
Convolutional neural networks (CNNs) are widely used for high-stakes applications like medicine, often surpassing human performance. However, most explanation methods rely on post-hoc attribution, approximating the decision-making process of already trained black-box models. These methods are often sensitive, unreliable, and fail to reflect true model reasoning, limiting their trustworthiness in critical applications. In this work, we introduce SoftCAM, a straightforward yet effective approach that makes standard CNN architectures inherently interpretable. By removing the global average pooling layer and replacing the fully connected classification layer with a convolution-based class evidence layer, SoftCAM preserves spatial information and produces explicit class activation maps that form the basis of the model’s predictions. Evaluated on three medical datasets spanning three imaging modalities, SoftCAM maintains classification performance while significantly improving both the qualitative and quantitative explanation compared to existing post-hoc methods. The code is available at https://github.com/kdjoumessi/SoftCAM.
keywords:
ElasticNet, Self-explainable models, Convolutional Neural Networks, Class Activation Maps, Attribution maps.
editors: Under Review for MIDL 2026
@@ 1 Introduction
Convolutional Neural Networks (CNNs) have revolutionized computer vision by effectively capturing local spatial patterns, reducing the number of parameters, and enabling faster convergence, leading to state-of-the-art performance in tasks such as image recognition and object detection [oquab2015object, li2021survey]. However, their lack of interpretability limits adoption in high-stakes fields such as medical image analysis, where transparency and trust are essential. To mitigate this issue, numerous post-hoc saliency- or attribution-based methods have been proposed to explain CNN predictions by highlighting “where” important features appear in the input through pixel-wise importance maps computed after training. Prominent approaches include class activation maps (CAM) techniques [zhou2016learning, selvaraju2017grad, jiang2021layercam], which combine convolutional features with class-specific weights; backpropagation-based methods [springenberg2014striving, sundararajan2017axiomatic], which propagate output gradients to the input to estimate pixel relevance; and perturbation-based methods [wang2020score, ivanovs2021perturbation], which systematically alter input features and measure the corresponding change in the model’s output.
While post-hoc attribution-based methods provide intuitive visualizations of discriminative regions, they frequently approximate rather than accurately reflect the model’s true reasoning [adebayo2018sanity, arun2021assessing, saporta2022benchmarking] and are often valid only under strong regularity assumptions [gunther2025informative]. Their post-hoc nature further limits their effectiveness—particularly in clinical applications [arun2021assessing]—due to low faithfulness, reliability, and consistency. Consequently, their visual outputs may not accurately reflect the model’s internal decision-making [adebayo2018sanity, saporta2022benchmarking]. Moreover, post-hoc techniques often struggle to precisely localize disease-relevant regions in medical images [arun2021assessing], where the scarcity of ground-truth annotations further complicates validation. To address these challenges, inherently or self-explainable models have been proposed [rudin2019stop], embedding interpretability directly into their architectures [brendel2018bagnets, chen2019looks, koh2020concept]. By coupling prediction with explanation, self-explainable models provide interpretable and transparent insights, in line with human reasoning. However, their reliance on specialized architectures limits generalization to widely used CNN models.
Motivated by these limitations, we introduce SoftCAM, a simple yet effective framework that makes CNNs inherently interpretable without relying on post-hoc explanation methods. SoftCAM generalizes the concept of class activation maps to transform conventional CNNs into self-explainable models. By removing the final global average pooling layer and replacing the fully connected classifier with a convolution-based class-evidence layer, SoftCAM converts standard CNNs into fully convolutional architectures that generate explicit, class-specific evidence maps used both for prediction and visual explanation. Evaluated on two widely used CNN architectures, we showed that the resulting SoftCAM-based models maintain competitive accuracy with their black-box baselines while achieving superior interpretability across three clinically relevant medical imaging datasets spanning three modalities. Furthermore, applying ElasticNet regularization, which combines ridge and lasso penalties, on evidence maps improves explanations both qualitatively and quantitatively, while revealing a task-dependent trade-off between sparsity and density. Finally, a comprehensive evaluation against five widely used post-hoc attribution methods showed that SoftCAM consistently outperforms these approaches on various explainability metrics.
@@ 2 Method
[…]
In contrast, traditional CNN are black-boxes: a global average pooling layer reduces $\mathbf{Z}$ to a vector of size $1\times D$, which is then fed into one or more linear fully connected layers (FCL) to generate the final prediction. Post-hoc methods are required to explain decisions.
@@ 2.1 CAM-based methods
Class Activation Maps (CAM) [zhou2016learning] are closely related to our approach, providing visual explanations of CNN predictions through class-specific saliency maps. CAM operates by linearly combining the final convolutional feature maps with their corresponding importance weights from the FCL classifier to produce class-wise attribution maps:
[…]
where $A_{k}(x_{1},x_{2})$ denotes the activation of the $k$ feature map at location $(x_{1},x_{2})$, and $w_{k}^{c}$ is the class-specific importance weight for the $k$ feature map from the fully connected layer.
Originally, CAM was designed for CNNs that use a GAP followed by a single FCL. Subsequent extensions, such as GradCAM [selvaraju2017grad] and LayerCAM [jiang2021layercam], introduced gradient-based approaches that use gradient to compute importance weights, enabling class-specific explanations for various architectures—particularly those with multiple FCLs after the GAP layer, such as VGG [simonyan2014very]. In GradCAM, the importance weights are computed by globally averaging the gradients of the target class score as $w_{k}^{c}=\frac{1}{N\times M}\sum_{i}^{N}\sum_{j}^{N}\frac{\partial y^{c}}{\partial A_{k}(i,j)}$, where $A_{k}(i,j)$ represents the activation at spatial location $(i,j)$ in the $k$-th feature map. Following GradCAM, several gradient-based and gradient-free extensions have emerged [he2022survey]. While gradient-based methods differ primarily in how they aggregate gradients to compute importance weights, gradient-free approaches estimate these weights without backpropagation, often relying on perturbation-based methods like ScoreCAM [wang2020score].
Despite their widespread clinical use [ayhan2022clinical], attribution-based methods have notable limitations [gunther2025informative]: they provide post-hoc explanations that may not reflect the model’s true reasoning (Appendix A). Gradient-based variants can suffer from gradient saturation and false confidence [wang2020score], while gradient-free approaches are computationally costly, requiring many forward passes on perturbed inputs.
2.2 Generalizing CAMs for self-explanability
Motivated by the limitations of post-hoc class activation map-based methods in interpreting CNN, we introduce SoftCAM (Fig. 1), a straightforward modification of black-box CNN classifiers that makes them inherently interpretable. SoftCAM achieves this by replacing the fully connected classification layer in classical CNNs with an explicit class-evidence convolutional layer, preserving spatial information and providing explanations in a single forward pass, eliminating the need and computational overhead for post-hoc techniques.
@@ 2.2 Generalizing CAMs for self-explanability
We make black-box CNN self-explainable by modifying how predictions are obtained. Any FCL of size $b_{1}\times b_{2}$ can be equivalently expressed as a $1\times 1$ convolutional layer with $b_{1}$ input channels and $b_{2}$ output channels. This allows systematic replacement of FCL classifier heads in CNN classifiers with convolutional layers, eliminating the GAP layer before classification while preserving model complexity and spatial localization. The new classifier module $h$ consists of convolutional layers (Fig. 1d) with $C$ convolution kernels of size $1\times 1$ and unit stride, producing class-specific evidence maps (Fig. 1e):
@@
where $\psi$ is a learnable parameter. These maps can be upsampled to the input resolution and overlaid on the image (Fig. 1g) for visualization and clinical interpretation. The module $h_{\psi}$ can be viewed as a soft, explainable generalization of classical post-hoc attribution methods (Eq. 1), mapping the high-dimensional feature map $\mathbf{Z}$ with $D$ channels into low-dimentional, class-wise activation maps $\mathbf{A}$ with $C$ channels corresponding to the target classes. Unlike CAM (Eq. 1) and related attribution methods which generate post-hoc heuristic explanations, our approach leverages the final feature map from the backbone and applies a parameterized function $h_{\psi}$ to directly produce class activation maps that are used for prediction. In contrast to classical CAM-based methods, the importance weights are not explicitly defined but are implicitly learned and encoded within the classifier’s parameters.
The resulting architecture is a fully convolutional, self-explainable model, where the predicted probabilities are derived directly from the evidence maps (Fig.1e), maintaining the model complexity without adding extra parameters:
[…]
Furthermore, the class evidence maps $\mathbf{A}$ serve as built-in explanations, directly representing the contribution of individual input regions to the final prediction (Fig. 1g). Using class-evidence maps for classification offers several advantages: all importance scores are weighted equally when computing class probabilities (Fig. 1f). Consequently, input feature patches with high activations in the evidence maps contribute more strongly to the prediction, analogous to linear models, where each input feature contributes linearly to the output.
@@ 2.3 Regularizing SoftCAM for interpretability
By using explicit class-evidence maps, SoftCAM-based models can be trained with regularization constraints directly applied to the explanation maps, enhancing interpretability. In practice, we apply an ElasticNet penalty [zou2005regularization], which linearly combines the $\ell_{1}$ (lasso) and $\ell_{2}$ (ridge) regularization, leading to the following loss function:
[…]
Here, CE denotes the cross-entropy loss, and $\mathbf{y}$ represents the reference labels. Setting $\lambda_{2}=0$, results in the lasso penalty promoting sparsity in evidence maps [donteu2023sparse] by suppressing less informative activations (mainly false positives), making it particularly useful for tasks where precision in explanations is crucial. In contrast, $\lambda_{1}=0$, gives the ridge penalty, which smooths activations without forcing them to zero, useful for assessing localization sensitivity over large regions, as it penalizes false negatives (Appendix E). ElasticNet thus provides a balance between lasso ($\lambda_{2}=0$) and ridge ($\lambda_{1}=0$) penalties.